### v1

# Hoeffding and Bernstein inequalities for weighted sums of exchangeable random variables

Abstract
The aim of this paper is to establish Hoeffding and Bernstein type concentration inequalities for weighted sums of exchangeable random variables. A special case is the i.i.d. setting, where random variables are sampled independently from some distribution (and are therefore exchangeable). In contrast to the existing literature on this problem, our results provide a natural unified view of both the i.i.d. and the exchangeable setting.
1 Introduction
Let $w_{1},\ldots,w_{n}\in\mathbb{R}$ be fixed weights and let $X_{1},\ldots,X_{n},X_{n+1},\ldots,X_{N}\in[-1,1]$ be exchangeable random variables. Our goal is to establish concentration properties for the weighted sum
$w_{1}X_{1}+\ldots+w_{n}X_{n}.$
In the special case where the $X_{i}$’s are i.i.d. (a strictly stronger condition), there are a wide range of well-known concentration bounds for this sum—for example, the Hoeffding and Bernstein inequalities, as well as multiple more technical results providing tighter bounds. There are also many recent results handling the exchangeable case—the $X_{i}$’s are no longer assumed to be independent but are required to satisfy exchangeability, meaning that the joint distribution of $(X_{1},\ldots,X_{N})$ is equal to that of any permutation, $(X_{\pi(1)},\ldots,X_{\pi(N)})$. In particular, tight bounds have been established for the special case where $w_{1}=\ldots=w_{n}$ (for example, if $w_{i}\equiv 1/n$, then the weighted sum can be viewed as a the sample mean, for a sample of size $n$ drawn without replacement from the finite population $X_{1},\ldots,X_{N}$).
However, for the general case where the $w_{i}$’s may be nonconstant, a key gap remains in the current literature on this problem, as we now describe. Since the i.i.d. assumption is a special case of exchangeability, we might expect that classical results for the i.i.d. setting (i.e., for the sum $w_{1}X_{1}+\ldots+w_{n}X_{n}$, where $X_{i}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$ for some distribution $P$) can be recovered as a special case of the concentration bounds for exchangeable sequences. Indeed, if we take $N\to\infty$, then an infinite exchangeable sequence $X_{1},X_{2},\ldots$ must be representable as an i.i.d. draw from some (random) distribution $P$, by De Finetti’s theorem (De Finetti, 1929; Hewitt and Savage, 1955); even if $N$ is finite, if $N\gg n$ then the joint distribution of an exchangeable sequence $X_{1},\ldots,X_{N}$ can be well approximated as an i.i.d. draw from some (random) $P$ (Diaconis and Freedman, 1980). Consequently, concentration results on the weighted sum $\sum_{i=1}^{n}w_{i}X_{i}$, under an assumption of exchangeability for $X_{1},\ldots,X_{N}$, should ideally recover the classical bounds for the i.i.d. case when we take $N\to\infty$. But, as we will see below, this is not the case for existing bounds when the weights $w_{1},\ldots,w_{n}$ are arbitrary.
The primary goal of this paper is to close this gap, by deriving concentration bounds that achieve the following aims.
Main aims: We seek to establish concentration results for the weighted sum $w_{1}X_{1}+\ldots+w_{n}X_{n}$ under the assumption of exchangeability, such that for a fixed number of terms $n$ in the sum, the results
  • (a)
    Provide (approximately) tight bounds for fixed $N$; and
  • (b)
    Recover classical bounds for the i.i.d. setting, in the limit where $N\to\infty$.
Moreover, since the distribution of the weighted sum $\sum_{i=1}^{n}w_{i}X_{i}$ remains the same if we permute the weights, for interpretability our results should
  • (c)
    Derive bounds that are symmetric in the weights $w_{1},\ldots,w_{n}$.
The results of this paper will establish both Hoeffding and Bernstein type bounds for the weighted sum in the exchangeable setting, which achieve these aims.
Notation.
For any integer $n\geq 1$, $[n]$ denotes the set $\{1,\ldots,n\}$. For a vector $x\in\mathbb{R}^{n}$, we define $\bar{x}=\frac{1}{n}\sum_{i=1}^{n}$ as its mean and $\sigma^{2}_{x}=\frac{1}{n}\sum_{i=1}^{n}(x_{i}-\bar{x})^{2}$ as its variance. For any $i\in[n]$, we will also write $x_{\geq i}$ to denote the subvector $(x_{i},\ldots,x_{n})\in\mathbb{R}^{n-i+1}$ (with $x_{\geq 1}=x$), and $x_{\leq i}$ for the subvector $(x_{1},\ldots,x_{i})\in\mathbb{R}^{i}$ (with $x_{\leq n}=x$).
Outline of paper.
We will begin by reviewing the classical Hoeffding and Bernstein bounds for the i.i.d. setting in Section 2. Our main results for the exchangeable setting will be presented in Section 3, where we will also discuss how the i.i.d. setting can be derived as a special case (as in aim (b) above). In Section 4 we will survey some related results in the literature, to compare our work to existing bounds. We will conclude with a short discussion in Section 5. All proofs are deferred to the Appendix.
2 Background
In this section, we review the well-known Hoeffding and Bernstein concentration inequalities for the i.i.d. setting, as well as some extensions to sampling without replacement.
2.1 Hoeffding’s inequality for the i.i.d. setting
Consider data points $X_{i}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, where $P$ is a distribution supported on $[-1,1]$ with mean $\mu_{P}$ and variance $\sigma^{2}_{P}$. As for the exchangeable case, we will study the distribution of the weighted sum $\sum_{i=1}^{n}w_{i}X_{i}$, for fixed weights $w_{1},\ldots,w_{n}$.
First we review the classical Hoeffding bound, which gives a concentration result for a sum of bounded independent terms. We begin with the Hoeffding bound on the moment-generating function (MGF) for the weighted sum $\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})$.
Theorem 1 (Hoeffding (1963)).
Let $w_{1},\ldots,w_{n}$ be fixed and let $X_{1},X_{2},\ldots\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, where $P$ is a distribution on $[-1,1]$ with mean $\mu_{P}$. For any $\lambda\in\mathbb{R}$ it holds that
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})\right\}}\right]\leq\exp\left\{\frac{\lambda^{2}\|w\|^{2}_{2}}{2}\right\}.$
In order to convert this MGF result to a bound on the tail probability, we will use the following fact (Boucheron et al., 2013, Section 2.3): for any random variable $Z\in\mathbb{R}$,
If $\mathbb{E}\left[{e^{\lambda Z}}\right]\leq e^{\lambda^{2}a^{2}/2}$ for all $\lambda\in\mathbb{R}$, then $\mathbb{P}\left\{{Z\geq a\sqrt{2\log(1/\delta)}}\right\}\leq\delta$ for all $\delta\in(0,1)$. (1)
Combining Theorem 1 with this fact immediately leads to the following corollary:
Corollary 2.
In the setting of Theorem 1, for any $\delta\in(0,1)$,
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})\geq\|w\|_{2}\sqrt{2\log(1/\delta)}}\right\}\leq\delta.$
2.2 Bernstein’s inequality for the i.i.d. setting
As above, we again consider the i.i.d. setting, $X_{i}\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$. In settings where the distribution $P$ has low variance, the Hoeffding bound can be extremely loose, since it characterizes the distribution of the weighted sum in terms of the maximum value of the $X_{i}$’s rather than the variance. In this regime, Bernstein’s inequality can give a substantially tighter bound. We begin with a bound on the MGF.
Theorem 3 (Bernstein (1946)).
Let $w_{1},\ldots,w_{n}$ be fixed and let $X_{1},X_{2},\ldots\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$, where $P$ is a distribution on $[-1,1]$ with mean $\mu_{P}$ and variance $\sigma^{2}_{P}$. For any $\lambda\in\mathbb{R}$ with $|\lambda|<\frac{3}{2\|w\|_{\infty}}$ it holds that
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})\right\}}\right]\leq\exp\left\{\frac{\lambda^{2}}{2(1-\frac{2\|w\|_{\infty}}{3}|\lambda|)}\cdot\sigma^{2}_{P}\|w\|^{2}_{2}\right\}.$
As for the Hoeffding bound, we can then convert this MGF bound to a tail probability bound by making use of the following fact (Boucheron et al., 2013, Section 2.4): for any random variable $Z\in\mathbb{R}$ and $a,b>0$,
[…]
Combining Theorem 3 with this fact immediately leads to the following corollary:
Corollary 4.
In the setting of Theorem 3, for any $\delta\in(0,1)$,
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})\geq\sigma_{P}\|w\|_{2}\sqrt{2\log(1/\delta)}+\frac{2}{3}\|w\|_{\infty}\log(1/\delta)}\right\}\leq\delta.$
For values of $\delta$ that are not too small, the first term $\sigma_{P}\|w\|_{2}\sqrt{2\log(1/\delta)}$ is usually dominant. Comparing to the Hoeffding bound in Corollary 2, then, the Bernstein bound differs in that (the dominant term in) the bound scales with the variance $\sigma_{P}$. This allows for a much tighter bound by using more information about the distribution $P$ than only its boundedness.
2.3 Do these bounds extend to exchangeable data?
Before turning to our main results, we first pause to ask whether these existing bounds hold also in the broader setting of exchangeable, rather than i.i.d., data. In fact, this question includes a very important special case: the unweighted sum, $\sum_{i=1}^{n}X_{i}$, or equivalently, weights $w_{1}=\ldots=w_{n}=1$. This problem has a long history in the literature, and is typically studied in the context of sampling without replacement from a finite population of size $N$. In fact, sampling without replacement has stronger concentration than an i.i.d. sample (i.e., sampling with replacement)—for instance, Serfling’s inequality (Serfling, 1974) establishes that
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}(X_{i}-\bar{X})\right\}}\right]\leq\exp\left\{\frac{n\lambda^{2}}{2}\cdot\left(1-\frac{n-1}{N}\right)\right\},$ (3)
which is strictly stronger than the Hoeffding bound for the i.i.d. case (Theorem 1) due to the additional factor $(1-\frac{n-1}{N})$. More generally, even if we only assume exchangeability of $X_{1},\ldots,X_{n}$, the Hoeffding and Bernstein bounds for the i.i.d. case still hold for the unweighted sum $\sum_{i=1}^{n}X_{i}$. (See Section 4.1 for more details on these results, and other results in this area.)
However, in the setting of general weights, this is no longer the case: as we will see in the following example, the bounds for the i.i.d. case may no longer hold.
Example 5.
Let $N=n=2$, and let $w_{1}=1$, $w_{2}=-1$, so that $\|w\|^{2}_{2}=2$. If $X_{1},X_{2}$ are sampled i.i.d. from the uniform distribution on $\{1,-1\}$, then we have
$\mathbb{E}\left[{e^{\lambda\sum_{i=1}^{2}w_{i}X_{i}}}\right]\leq e^{\frac{\lambda^{2}\|w\|^{2}_{2}}{2}}$
by the Hoeffding bound (Theorem 1). If instead $X_{1},X_{2}$ are sampled uniformly without replacement from $\{1,-1\}$—that is, $(X_{1},X_{2})$ is equal to either $(-1,1)$ or $(1,-1)$, each with probability $0.5$—then we have
$\mathbb{E}\left[{e^{\lambda\sum_{i=1}^{2}w_{i}X_{i}}}\right]=\mathbb{E}\left[{e^{\lambda(X_{1}-X_{2})}}\right]=\frac{e^{2\lambda}+e^{-2\lambda}}{2}\approx e^{2\lambda^{2}}=e^{2\cdot\frac{\lambda^{2}\|w\|^{2}_{2}}{2}},$
where the approximation holds for $\lambda\approx 0$. That is, for this exchangeable distribution on $X_{1},X_{2}$, the Hoeffding bound for the i.i.d. case (Theorem 1) no longer holds.
To summarize, we have seen that the unweighted case and the weighted case are fundamentally quite different:
  • In the unweighted setting, the sum $\sum_{i=1}^{n}X_{i}$ (when $X_{i}$’s are sampled without replacement from a finite population) concentrates more strongly than in the i.i.d. case; bounds for the i.i.d. setting will automatically apply to this problem. Therefore, the main focus of the existing literature on this problem is to establish bounds for sampling without replacement that are tighter than the i.i.d. bounds (such as Serfling’s inequality (3), which is tighter than the i.i.d. Hoeffding bound).
  • In contrast, in the weighted setting, as demonstrated by Example 5, if the $X_{i}$’s are sampled without replacement (or, more generally, are exchangeable), for nonconstant weights the sum $\sum_{i=1}^{n}w_{i}X_{i}$ may not satisfy the i.i.d. bounds. Our goal is therefore to derive bounds that are as close as possible to the i.i.d. bounds.
With this motivation in place, we are ready to turn to our main results.
3 Main results
We now present our new bounds for the case of exchangeable data: a Hoeffding bound in Section 3.1, and a Bernstein bound in Section 3.2. We will then see in Section 3.3 how these results lead to the classical i.i.d. bounds as a special case.
3.1 A Hoeffding bound for the exchangeable case
Our first result is a Hoeffding type bound. First we need a definition: for any $n\geq 2$ we define
$\epsilon_{n}=\frac{H_{n}-1}{n-H_{n}}\textnormal{ where $H_{n}=1+\frac{1}{2}+\ldots+\frac{1}{n}$}.$
Since the harmonic series scales as $H_{n}\asymp\log n$, this means that $\epsilon_{n}=\mathcal{O}(\frac{\log n}{n})$.
Theorem 6.
Let $w\in\mathbb{R}^{n}$ be fixed and let $X_{1},\ldots,X_{N}\in[-1,1]$ be exchangeable, where $N\geq n\geq 2$. Then for any $\lambda\in\mathbb{R}$,
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right\}}\right]\leq\exp\left\{\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\cdot(1+\epsilon_{N})\right\}.$
Next we will see an important implication of this result. In the i.i.d. case, the Hoeffding MGF bound leads to a subgaussian tail bound on the weighted sum. The following corollary, which is an immediate consequence of Theorem 6 together with the fact (1), shows an analogous result for the exchangeable case.
Corollary 7.
In the setting of Theorem 6, for any $\delta\in(0,1)$,
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\geq\|w\|_{2}\cdot\sqrt{2(1+\epsilon_{N})\cdot\log(1/\delta)}}\right\}\leq\delta.$
Comparing to the i.i.d. case studied in Theorem 1 and Corollary 2, aside from replacing the mean $\mu_{P}$ with an empirical mean $\bar{X}$, we see one difference: the term $(1+\epsilon_{N})$ that appears in the bounds for our two new results. Importantly, this term involves $\epsilon_{N}$ (which vanishes with $N$, the length of the exchangeable sequence), rather than $\epsilon_{n}$ (which may be relatively large if the number of terms $n$ in the weighted sum is small).
3.1.1 A special case: nonnegative weights
Recall from Section 2.3 that the the i.i.d. Hoeffding bound (Theorem 1) holds for the unweighted case, $w_{1}=\ldots=w_{n}=1$, under the broader assumption of exchangeability. While Example 5 demonstrates that this is no longer true in general for arbitrary weights, we will now see that nonnegative weights avoid this issue.
Theorem 8.
Let $w\in\mathbb{R}^{n}$ be fixed and let $X_{1},\ldots,X_{n}\in[-1,1]$ be exchangeable. If $w_{i}\geq 0$ for all $i$, then for any $\lambda\in\mathbb{R}$,
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right\}}\right]\leq\exp\left\{\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\right\}.$
Consequently, by (1), we obtain a tail bound that is as strong as for the i.i.d. case.
Corollary 9.
In the setting of Theorem 8, for any $\delta\in(0,1)$,
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\geq\|w\|_{2}\cdot\sqrt{2\log(1/\delta)}}\right\}\leq\delta.$
3.2 A Bernstein bound for the exchangeable case
We next present a Bernstein type bound, which offers tighter control of the MGF in settings where $\sigma^{2}_{X}$ is small.
Theorem 10.
Let $w\in\mathbb{R}^{n}$ be fixed and let $X_{1},\ldots,X_{N}\in[-1,1]$ be exchangeable, where $N\geq n\geq 2$. Then for any $\lambda\in\mathbb{R}$ with $|\lambda|<\frac{3}{2\|w\|_{\infty}(1+\epsilon_{N})}$,
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})-\frac{\lambda^{2}(1+\epsilon_{N})}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty}(1+\epsilon_{N}))}\cdot\tilde{\sigma}^{2}_{X,N}\|w\|^{2}_{2}\right\}}\right]\leq 1,$
where $\tilde{\sigma}_{X,N}^{2}=\sigma^{2}_{X}+4\epsilon_{N}$.
As for the i.i.d. case, this MGF bound leads to a tail bound, by combining Theorem 10 with the fact (2) (with a few extra steps to account for the fact that $\tilde{\sigma}^{2}_{X,N}$ is random rather than fixed).
Corollary 11.
In the setting of Theorem 10, for any $\delta\in(0,1)$,
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\geq\tilde{\sigma}_{X,N}\|w\|_{2}\sqrt{2(1+\epsilon_{N})\log(1/\delta)}+\frac{2}{3}\|w\|_{\infty}(1+\epsilon_{N})\log(1/\delta)}\right\}\leq\delta.$
To compare to the results given in Theorem 3 and Corollary 4 for the i.i.d. case, these new results have the following differences. The terms $\mu_{P}$ and $\sigma^{2}_{P}$ are replaced with the empirical mean of the vector, $\bar{X}$, and with an adjusted empirical variance $\tilde{\sigma}^{2}_{X,N}$ (which is only slightly larger than $\sigma^{2}_{X}$, if $N$ is large). In addition, as for the Hoeffding results, an inflation term $(1+\epsilon_{N})$ appears in several places in these bounds. Again, we note that the effect of this inflation term is vanishing with $N$ rather than with the (potentially smaller value) $n$.
3.3 From exchangeable to i.i.d.
Next, we will show that the original i.i.d. versions of the Hoeffding and Bernstein bounds, given in Theorems 1 and 3, can be recovered from our new results for exchangeable data, by taking $N\to\infty$. This is the goal that we posed in aim (b) of Section 1.
Consider an infinite sequence $X_{1},X_{2},\ldots\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$. Recall that $X_{\leq N}=(X_{1},\ldots,X_{N})$ denotes the vector containing the first $N$ draws. Note that $X_{\leq N}$ is exchangeable, since i.i.d. is a special case of exchangeability. By the strong law of large numbers, we have $\bar{X}_{\leq N}\to\mu_{P}$ and $\sigma^{2}_{X_{\leq N}}\to\sigma^{2}_{P}$ almost surely. We will now see that the convergence of these quantities allows us to recover the i.i.d. results, Theorems 1 and 3, as special cases of our new bounds.
Specifically, for the Hoeffding bound, since $\bar{X}_{\leq N}\to\mu_{P}$ almost surely (and also $\epsilon_{N}\to 0$), for any $\lambda\in\mathbb{R}$ we have
$\lim_{N\to\infty}\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X}_{\leq N})-\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\cdot(1+\epsilon_{N})\right\}=\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})-\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\right\},$
almost surely. Therefore, by the dominated convergence theorem,
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})-\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\right\}}\right]\\ =\lim_{N\to\infty}\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X}_{\leq N})-\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\cdot(1+\epsilon_{N})\right\}}\right]\leq 1,$
where for the last step we apply our new result, Theorem 6. This is equivalent to the bound given in Theorem 1 for the i.i.d. case.
Next we turn to the Bernstein bound. Since $\bar{X}_{\leq N}\to\mu_{P}$ and $\sigma^{2}_{X_{\leq N}}\to\sigma^{2}_{P}$ almost surely (and also $\epsilon_{N}\to 0$, which additionally implies $\tilde{\sigma}^{2}_{X_{\leq N},N}\to\sigma^{2}_{P}$ almost surely), we then have
$\lim_{N\to\infty}\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X}_{\leq N})-\frac{\lambda^{2}(1+\epsilon_{N})}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty}(1+\epsilon_{N}))}\tilde{\sigma}^{2}_{X_{\leq N},N}\|w\|^{2}_{2}\right\}\\ =\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\sigma^{2}_{P}\|w\|^{2}_{2}\right\},$
almost surely for any $\lambda\in\mathbb{R}$. Now assume $|\lambda|<\frac{2}{3\|w\|_{\infty}}$. Note that, since $\epsilon_{N}\to 0$, for sufficiently large $N$ we have $|\lambda|<\frac{2}{3\|w\|_{\infty}(1+\epsilon_{N})}$. Following the same arguments as for the Hoeffding case above, by applying Theorem 10 to each exchangeable sequence $X_{\leq N}$, along with the dominated convergence theorem, we obtain the bound
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\mu_{P})-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\sigma^{2}_{P}\|w\|^{2}_{2}\right\}}\right]\\ \leq 1,$
which recovers Theorem 3’s bound for the i.i.d. case.
4 Related work
We now compare our new Hoeffding and Bernstein bounds to related results in the literature.
4.1 Existing results for special cases
We first summarize some existing results for special cases of the problem studied here.
The i.i.d. setting.
First, as discussed earlier, the i.i.d. setting, where $X_{1},X_{2},\ldots\stackrel{{\scriptstyle\textnormal{iid}}}{{\sim}}P$ for some distribution $P$, is of course a special case of exchangeability (and can be viewed as taking a limit $N\to\infty$ in the exchangeable setting). In addition to the classical Hoeffding bound (Hoeffding, 1963) and Bernstein bound (Bernstein, 1946) for the i.i.d. setting (presented in Theorems 1 and 3), many other types of concentration bounds are well established in the literature, and may provide sharper bounds for certain settings. See Boucheron et al. (2013) for background on a range of such bounds. More recent literature has considered the extension to time-uniform bounds—providing bounds that hold uniformly over a sequence of times $n$ rather than at a single fixed $n$. In particular, the work of Howard et al. (2020, 2021) provides a martingale based approach to this problem (this is closely related to time-uniform bounds for sampling without replacement, studied by Waudby-Smith and Ramdas (2020), whose work we describe in more detail below).
The unweighted sum setting.
Another special case is that of an unweighted sum, $\sum_{i=1}^{n}X_{i}$ (i.e., we take weights $w_{i}\equiv 1$), as we described briefly in Section 2.3 above. Hoeffding (1963, Theorem 4) establishes an important result for this setting: for any continuous and convex function $f$,
$\mathbb{E}\left[{f(X_{1}+\ldots+X_{n})}\right]\leq\mathbb{E}\left[{f(X^{\prime}_{1}+\ldots+X^{\prime}_{n})}\right],$ (4)
where $X_{1},\ldots,X_{n}$ are sampled uniformly without replacement from a finite population of size $N$, while $X^{\prime}_{1},\ldots,X^{\prime}_{n}$ are sampled uniformly with replacement from this same population (and therefore, are i.i.d.). In particular, taking $f(t)=e^{\lambda t}$, this implies that the Hoeffding and Bernstein bounds for the i.i.d. setting, are also valid for sampling without replacement (or, more generally, exchangeable data), when we take weights $w_{i}\equiv 1$.
As we discussed earlier in Section 2.3 (see (3)), Hoeffding’s inequality can be strengthened further to Serfling’s inequality (Serfling, 1974) for the setting of sampling without replacement. Multiple works in the more recent literature have obtained additional refined bounds for this problem. For example, the work of Bardenet and Maillard (2015) gives a tighter version of the Hoeffding–Serfling inequality, and also derives some Bernstein type bounds. Stronger Bernstein type results are established by Greene and Wellner (2017), via relating the distribution of the unweighted sum $\sum_{i=1}^{n}X_{i}$ to a Hypergeometric distribution (which can be viewed as a special case—i.e., sampling without replacement from a finite population that contains only 00’s and $1$’s).
While these results consider a problem that is clearly closely related to the central question of this work, the general case (where weights $w_{i}$ may not be constant) is a more challenging problem. As we saw in Example 5, the unweighted case is fundamentally different from the general case, where the $w_{i}$’s may vary. In general, this means that the techniques from this part of the literature are not typically possible to extend to the broader setting of a weighted sum.
4.2 Existing results for weighted sums
Next, we compare to existing results for the more general problem considered in this paper, i.e., the weighted sum $\sum_{i=1}^{n}w_{i}X_{i}$, for arbitrary fixed weights $w_{i}$ (and exchangeable $X_{i}$’s).
First, we present the result of Gan et al. (2010, Theorem 1), which proves a Hoeffding type bound using an elegant argument based on Stein’s exchangeable pairs. Translating their result into our notation, their theorem implies that
$\mathbb{P}\left\{{\left|\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right|\geq\|w\|_{2}\sqrt{4\log(2/\delta)}}\right\}\leq\delta,$
for any $\delta\in(0,1)$. Comparing this to our result in Corollary 7, our results instead yield the tighter bound
$\mathbb{P}\left\{{\left|\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right|\geq\|w\|_{2}\sqrt{2(1+\epsilon_{N})\cdot\log(2/\delta)}}\right\}\leq\delta.$
(Note that the term $\log(1/\delta)$, appearing in the statement of the corollary, has been replaced with $\log(2/\delta)$—this is because we are now bounding the absolute deviation, $|\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})|$, rather than a one-sided bound.) Since $\epsilon_{N}\leq 1$ for all $N\geq 2$, and $\epsilon_{N}\to 0$ as $N\to\infty$, our bound is strictly tighter (in terms of the constant) as compared to the result of Gan et al. (2010) for any $N>2$. In particular, only our bound achieves the same constant as the i.i.d. bound in the limit $N\to\infty$.
We next turn to the recent results of Waudby-Smith and Ramdas (2020). The central problem studied in this work is different from our work. If we observe $X_{1},\ldots,X_{n}$ as a sample drawn without replacement from a finite population of size $N$, our aim is to bound deviations of $\sum_{i=1}^{n}w_{i}X_{i}$ for fixed weights $w_{i}$. In particular, for nonconstant weights, this question is nontrivial even if $\bar{X}$, the mean of the finite population, is known—or even if all the values $\{X_{1},\ldots,X_{N}\}$ in the finite population are known. In contrast, Waudby-Smith and Ramdas (2020)’s aim is to estimate the unknown mean $\bar{X}$ using the sample $X_{1},\ldots,X_{n}$. They do this by designing weights $w_{1},\ldots,w_{n}$ such that the sum, $\sum_{i=1}^{n}w_{i}X_{i}$, can be used to construct a tight confidence intervals for $\bar{X}$. Despite these differences, these two questions are closely related, since a bound of the form $\left|\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right|\leq\Delta_{n}$ for some bound $\Delta_{n}$ (as we provide in our main results here) can equivalently be viewed as a confidence interval for the mean $\bar{X}$, namely, $\bar{X}\in\frac{\sum_{i=1}^{n}w_{i}X_{i}}{\sum_{i=1}^{n}w_{i}}\pm\Delta_{n}$.
Another key distinction is that Waudby-Smith and Ramdas (2020) study the streaming setting, providing a time-uniform guarantee (i.e., a valid confidence interval for $\bar{X}$ simultaneously at every $n=1,\ldots,N$), which is a stronger guarantee than the fixed-$n$ case considered in this work. As a consequence, though, the problem they study is not invariant to permuting indices, due to the sequential nature of the streaming setting. In contrast, as described in aim (c) in Section 1, since we are only considering a fixed value of $n$ in our work, any interpretable bound in our setting should be invariant to reordering the indices, and therefore our results are not easily comparable with theirs. However, the results of Waudby-Smith and Ramdas (2020) provide some key technical tools that we use in this work for proving our main results. We will discuss this further, and give more details on the bounds derived by Waudby-Smith and Ramdas (2020), in Appendix B. We remark also that Shekhar et al. (2023, Appendix D) extend these results to allow for a setting where the $X_{i}$’s are sampled with a weighted, rather than uniform, distribution.
The matrix setting.
A more challenging problem that has also been considered in the literature is the following: given a fixed matrix $A\in\mathbb{R}^{n\times n}$, derive concentration bounds for the sum
$\sum_{i=1}^{n}A_{i,\pi(i)},$
where the permutation $\pi\in\mathcal{S}_{n}$ is sampled uniformly at random (here $\mathcal{S}_{n}$ denotes the set of permutations of $[n]$). To connect this to the problem of bounding a weighted sum $\sum_{i=1}^{n}w_{i}X_{i}$, we consider a fixed vector $x=(x_{1},\ldots,x_{n})\in[-1,1]^{n}$, and let $X_{i}=x_{\pi(i)}$ be defined by a random permutation of $x$ (equivalently, by sampling without replacement from the entries of $x$). If we then consider the rank-$1$ matrix $A=wx^{\top}$, we have $\sum_{i=1}^{n}w_{i}X_{i}=\sum_{i=1}^{n}A_{i,\pi(i)}$. In other words, any result bounding the sum in the matrix setting, can be applied to the setting of a weighted sum of exchangeable random variables as a special case.
While many results of this type have been shown, here we mention several recent results that comprise some of the strongest known bounds. If we assume $A_{ij}\in[0,1]^{n\times n}$, then Chatterjee (2007) proves that
$\mathbb{P}\left\{{\left|\sum_{i=1}^{n}A_{i,\pi(i)}-\mathbb{E}\left[{\sum_{i=1}^{n}A_{i,\pi(i)}}\right]\right|\geq t}\right\}\leq 2\exp\left\{-\frac{t^{2}}{\frac{4}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}+2t}\right\},$
while Albert (2019) proves that
$\mathbb{P}\left\{{\left|\sum_{i=1}^{n}A_{i,\pi(i)}-\mathbb{E}\left[{\sum_{i=1}^{n}A_{i,\pi(i)}}\right]\right|\geq t}\right\}\leq 8e^{1/16}\exp\left\{-\frac{t^{2}}{16\left(\frac{4}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}^{2}+2t\right)}\right\}.$
While the latter result in particular appears to be a Bernstein type bound, neither one of these bounds implies results of the form of our theorems. In particular, the bounds above are favorable only if the values of $A$ concentrate near zero, so that $\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}$ or $\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}^{2}$ is small; this does not translate to a useful result for our setting, where the $w_{i}$’s and $X_{i}$’s may be positive or negative.
The more recent work of Polaczyk (2023) offers a result that is more directly comparable to our work. Their result considers a matrix $A\in[-1,1]^{n\times n}$, and assumes that the matrix is centered, $\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}=0$. They prove
$\mathbb{P}\left\{{\sum_{i=1}^{n}A_{i,\pi(i)}\geq t}\right\}\leq 2\exp\left\{-\frac{t}{36}\cdot\log\left(1+\frac{t}{\frac{36}{n}\sum_{i=1}^{n}\sum_{j=1}^{n}A_{ij}^{2}}\right)\right\}.$
To compare this to our work, suppose $\|w\|_{\infty}\leq 1$, and assume again that $X_{i}=x_{\pi(i)}$ for a fixed vector $x_{1},\ldots,x_{n}$. Consider the simple case $\bar{x}=0$. Then $A=wx^{\top}$ is centered and takes values in $[-1,1]$, and so Polaczyk (2023)’s bound can be applied to yield
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\geq t}\right\}\leq 2\exp\left\{-\frac{t}{36}\cdot\log\left(1+\frac{t}{36\sigma^{2}_{X}\|w\|^{2}_{2}}\right)\right\}.$
Since we have $\log(1+\epsilon)\geq\frac{\epsilon}{1+\epsilon}$ for any $\epsilon>0$ (and this inequality is an approximate equality if $\epsilon\approx 0$), we can relax this to the bound
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\geq t}\right\}\leq 2\exp\left\{-\frac{t^{2}}{36\left(36\sigma^{2}_{X}\|w\|^{2}_{2}+t\right)}\right\},$
which is nearly the same in the regime $\frac{t}{36\sigma^{2}_{X}\|w\|^{2}_{2}}\approx 0$. This implies
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\geq 36\sigma_{X}\|w\|_{2}\sqrt{\log(2/\delta)}+36\log(2/\delta)}\right\}\leq\delta$
for any $\delta\in(0,1)$. Up to constants, this bound has the same form as our Bernstein result for the exchangeable setting (see Corollary 11). However, the constants are substantially larger, and so this bound is comparatively very loose (and will not recover the classical bound for the i.i.d. case).
4.3 Additional related work
Finally, we mention a related line of work, which studies the problem of establishing a central limit theorem for sums of exchangeable, rather than i.i.d. (or independent), random variables. Hájek (1960) establishes a central limit theorem for the unweighted case, $\sum_{i=1}^{n}X_{i}$ (where $X_{i}$’s are sampled without replacement from a finite population). Bolthausen (1984) studies the problem for the matrix case, establishing a central limit theorem for the sum $\sum_{i=1}^{n}A_{i,\pi(i)}$ (which, as discussed above, can be specialized to the problem of a weighted sum, $\sum_{i=1}^{n}w_{i}X_{i}$, for exchangeable $X_{1},\ldots,X_{n}$). More recent work by Li and Ding (2017) considers a different extension, which is important for the causal inference literature: they establish a multivariate central limit theorem for the random vector $(\bar{X}^{(1)},\ldots,\bar{X}^{(G)})$, where a finite population with values $x_{1},\ldots,x_{N}$ is randomly partitioned into $G$ groups, and $\bar{X}^{(g)}$ denotes the sample mean in group $g$ for each $g\in[G]$. (Note that each individual term, $\bar{X}^{(g)}$, can be viewed as a sum of random variables sampled without replacement—namely, sampling the values assigned to group $g$.)
5 Discussion
We conclude with a brief summary of our main results, and an exploration of some questions raised by this work.
Summary.
In this work, we have established new versions of Hoeffding and Bernstein inequalities for weighted sums of exchangeable random variables, $w_{1}X_{1}+\ldots+w_{n}X_{n}$ for an exchangeable sequence $X_{1},\ldots,X_{n},X_{n+1},\ldots,X_{N}$. Compared to prior work, these new bounds offer the benefit of reducing to existing results for the i.i.d. setting in the limit $N\to\infty$. Essentially, we can think of these results as providing a continuous interpolation between the setting of a random permutation (where $N=n$, and so $X_{1},\ldots,X_{n}$ is simply a random permutation of a finite list) and the setting of i.i.d. sampling (where $N=\infty$).
Open questions.
Examining these results in the context of the existing literature raises a number of open questions and potential extensions. First, it is natural to ask whether the bounds are tight: for instance, are the inflation terms $(1+\epsilon_{N})$ unavoidable? We have seen that in the special case of nonnegative weights $w_{i}\geq 0$, this inflation can be avoided for the Hoeffding bound (and the i.i.d. Hoeffding bound is recovered). However, it is unclear if the same might be true for the Bernstein bound with nonnegative weights. In addition, in the simplest setting where $w_{1}=\ldots=w_{n}=1$, Serfling’s inequality (3) gives a result that is strictly stronger than in the i.i.d. case—is there a more general result for nonnegative weights that can similarly capture improved concentration relative to the i.i.d. setting?
Next, for the i.i.d. setting, recent lines of work have established concentration bounds that are stronger or more flexible than the Hoeffding and Bernstein bounds—for example, empirical Bernstein inequalities, where the variance $\sigma^{2}_{P}$ of the distribution can be replaced by an empirical estimate (see Boucheron et al. (2013) for a survey of these types of results). Waudby-Smith and Ramdas (2020)’s recent work provides empirical Bernstein bounds for exchangeable data in the streaming setting. Can empirical Bernstein type bounds (or other types of bounds) be derived for exchangeable data in a way that satisfies the aims (a), (b), and (c) introduced in Section 1?
Finally, in Section 4 we also presented some existing results on a matrix version of the problem, where the sum is given by $\sum_{i=1}^{n}A_{i,\pi(i)}$ for a fixed matrix $A\in\mathbb{R}^{n\times n}$ and a random permutation $\pi$ (recall that the weighted sum, $\sum_{i=1}^{n}w_{i}X_{i}$, can be reinterpreted as a version of this problem where $A=wX^{\top}$ has rank $1$). It is possible that the technical tools developed in this work could be extended to derive new bounds for this matrix setting as well.
Acknowledgements
The author was supported by the Office of Naval Research via grant N00014-20-1-2337 and by the National Science Foundation via grant DMS-2023109. The author thanks Aaditya Ramdas and Ian Waudby-Smith for helpful discussions.
Appendix A Background: proofs for the i.i.d. case
Before presenting proofs of our new results, we begin by reviewing the standard proofs of the classical Hoeffding and Bernstein bounds, to build intuition and introduce some useful inequalities.
Proof of Theorem 1.
The basic Hoeffding inequality states that, if $Z$ is a mean-zero random variable with $a\leq Z\leq b$ almost surely, then
$\mathbb{E}\left[{e^{tZ}}\right]\leq e^{t^{2}(b-a)^{2}/8}.$ (5)
For each $i\in[n]$, we have $-1-\mu_{P}\leq X_{i}-\mu_{P}\leq 1-\mu_{P}$ almost surely, and so applying (5) with $Z=X_{i}-\mu_{i}$, $t=\lambda w_{i}$, $a=-1-\mu_{P}$, and $b=1-\mu_{P}$, we have
$\mathbb{E}\left[{e^{\lambda w_{i}(X_{i}-\mu_{P})}}\right]\leq e^{\lambda^{2}w_{i}^{2}/2}.$
Taking the product over all $i$ completes the proof. ∎
Proof of Theorem 3.
The basic Bernstein inequality states that, if $Z$ is a mean-zero random variable with $\textnormal{Var}(Z)=\sigma^{2}$ and $|Z|\leq B$ almost surely, then
$\mathbb{E}\left[{e^{tZ}}\right]\leq e^{\frac{t^{2}\sigma^{2}}{2(1-|t|B/3)}}$ (6)
for any $t\in\mathbb{R}$ with $|t|<3/B$. For each $i\in[n]$, we have $|X_{i}-\mu_{P}|\leq|X_{i}|+|\mu_{P}|\leq 2$ almost surely, and so applying (6) with $Z=X_{i}-\mu_{P}$, $t=\lambda w_{i}$, $\sigma^{2}=\sigma^{2}_{P}$, and $B=2$, we obtain
$\mathbb{E}\left[{e^{\lambda w_{i}(X_{i}-\mu_{P})}}\right]\leq\exp\left\{\frac{\lambda^{2}w_{i}^{2}\sigma^{2}_{P}}{2\left(1-\frac{2|\lambda|\|w\|_{\infty}}{3}\right)}\right\}.$
Taking the product over all $i$ establishes the bound. ∎
Appendix B Preliminaries: martingale based results
In this section, we present MGF bounds for the weighted sum that are based on martingale arguments—for the Hoeffding case, we will review an existing bound from the work of Waudby-Smith and Ramdas [2020], while for the Bernstein case, we will present a new bound that is similar in spirit to their work. As we will see shortly, these bounds cannot be immediately applied to establish the type of results that are the aim of this paper, but they form an intermediate step towards proving our theorems.
B.1 A martingale result for the Hoeffding bound
The following result is derived by Waudby-Smith and Ramdas [2020]. We present it here with different notation to maintain consistency within our paper.
Proposition 12 (Waudby-Smith and Ramdas [2020, Theorem 3.1]).
Fix $N\geq n\geq 1$ and $w\in\mathbb{R}^{n}$, and let $X_{1},\ldots,X_{N}\in[-1,1]$ be exchangeable. Fix any $\lambda\in\mathbb{R}$, and define $M_{0}=1$ and
$M_{k}=\exp\left\{\lambda\sum_{i=1}^{k}w_{i}(X_{i}-\bar{X}_{\geq i})-\frac{\lambda^{2}}{2}\sum_{i=1}^{k}w_{i}^{2}\right\}$
for $k\in[n]$. Then $M_{k}$ is a supermartingale, i.e., $\mathbb{E}\left[{M_{k}}\ \middle|\ {M_{k-1}}\right]\leq M_{k-1}$ for all $k\in[n]$.
Applying this proposition, we have $\mathbb{E}\left[{M_{n}}\right]\leq\mathbb{E}\left[{M_{0}}\right]=1$, and so
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X}_{\geq i})-\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\right\}}\right]\leq 1.$
This can be used to prove a tail bound on the weighted sum $\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})$—namely, by (1), we have
$\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\leq\sum_{i=1}^{n}w_{i}(\bar{X}_{\geq i}-\bar{X})+\|w\|_{2}\sqrt{2\log(1/\delta)}$
with probability $\geq 1-\delta$.
However, the term $\sum_{i=1}^{n}w_{i}(\bar{X}_{\geq i}-\bar{X})$ appearing in the upper bound is unsatisfactory, for two reasons: firstly, the term is not invariant to the ordering of the indices (as in aim (c)), and second, the term is random and therefore would itself need to be bounded. If $N-n$ is sufficiently large, then this is certainly possible, since for each $i\in[n]$, we can establish concentration properties for $\bar{X}_{\geq i}-\bar{X}$. But in the case $N=n$, say, the term may be quite large, meaning that we have not achieved sharp bounds for fixed $N$ (aim (a)). Nonetheless, we will see later on how this martingale-based MGF bound can be used to prove Theorem 6.
B.2 A martingale result for the Bernstein bound
Next, we present a Bernstein type version of this result. The following proposition is new, but this result builds on recent techniques and related bounds from the literature, e.g., Waudby-Smith and Ramdas [2020, Theorem 3.2], Howard et al. [2020, Corollary 1(c)], Waudby-Smith and Ramdas [2024, Section 5].
Proposition 13.
Fix $n\geq 1$ and $w\in\mathbb{R}^{n}$, and let $X=(X_{1},\ldots,X_{n})$ where $X_{1},\ldots,X_{n}\in[-1,1]$ are exchangeable. Fix any $\lambda\in\mathbb{R}$ satisfying $|\lambda|<\frac{2}{3\|w\|_{\infty}}$, and define $M_{0}=1$, and
$M_{k}=\exp\left\{\lambda\sum_{i=1}^{k}w_{i}(X_{i}-\bar{X}_{\geq i})-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\sum_{i=1}^{k}w_{i}^{2}\sigma^{2}_{X_{\geq i}}\right\}$
for $k\in[n]$. Then $M_{k}$ is a supermartingale, i.e., $\mathbb{E}\left[{M_{k}}\ \middle|\ {M_{k-1}}\right]\leq M_{k-1}$ for all $k\in[n]$.
As for the Hoeffding case, let us now examine the implications of this result. Applying the proposition, we have
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X}_{\geq i})-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\sum_{i=1}^{n}w_{i}^{2}\sigma^{2}_{X_{\geq i}}\right\}}\right]\leq 1.$
This leads to a Bernstein tail type bound on the weighted sum $\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})$: applying (2) yields
$\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\leq\sum_{i=1}^{n}w_{i}(\bar{X}_{\geq i}-\bar{X})+\sqrt{\sum_{i=1}^{n}w_{i}^{2}\sigma^{2}_{X_{\geq i}}}\sqrt{2\log(1/\delta)}+\frac{2}{3}\|w\|_{\infty}\log(1/\delta)$
with probability $\geq 1-\delta$. But as for the Hoeffding setting, this bound is again unsatisfactory, because there are terms in the upper bound that are not invariant to the ordering of the indices (violating aim (c)) and may be highly variable when $N=n$ (violating aim (a)). We will see later on, however, that this martingale-based MGF bound can be used to prove Theorem 10.
Proof of Proposition 13.
By construction, for each $k\in[n]$ we can write
$M_{k}=M_{k-1}\cdot\exp\left\{\lambda w_{k}(X_{k}-\bar{X}_{\geq k})-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}w_{k}^{2}\sigma^{2}_{X_{\geq k}}\right\}.$
Define a filtration $\{\mathcal{F}_{k}\}_{k\in\{0,\ldots,n\}}$, where $\mathcal{F}_{k}$ is the $\sigma$-algebra generated by $\bar{X},\sigma^{2}_{X},X_{1},\ldots,X_{k}$. For any $i\leq k$, we can calculate
$\bar{X}_{\geq i}=\frac{X_{i}+\ldots+X_{n}}{n-i+1}=\frac{n\bar{X}-(X_{1}+\ldots+X_{i-1})}{n-i+1},$
and so $\bar{X}_{\geq i}$ is $\mathcal{F}_{k-1}$-measurable. Furthermore,
$\sigma^{2}_{X_{\geq i}}=\frac{1}{n-i+1}\sum_{j=i}^{n}(X_{j}-\bar{X}_{\geq i})^{2}=\frac{1}{n-i+1}\sum_{j=i}^{n}(X_{j}-\bar{X})^{2}-(\bar{X}_{\geq i}-\bar{X})^{2}\\ =\frac{n\sigma^{2}_{X}}{n-i+1}-\frac{1}{n-i+1}\sum_{j=1}^{i-1}(X_{j}-\bar{X})^{2}-(\bar{X}_{\geq i}-\bar{X})^{2}.$
Since each of these three terms is $\mathcal{F}_{k-1}$-measurable, this means that $\sigma^{2}_{X_{\geq i}}$ is $\mathcal{F}_{k-1}$-measurable, as well. This also implies $M_{k}$ is $\mathcal{F}_{k}$-measurable for each $k$. We then have
$\mathbb{E}\left[{M_{k}}\ \middle|\ {\mathcal{F}_{k-1}}\right]=M_{k-1}\cdot\exp\left\{-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}w_{k}^{2}\sigma^{2}_{X_{\geq k}}\right\}\cdot\mathbb{E}\left[{\exp\left\{\lambda w_{k}(X_{k}-\bar{X}_{\geq k})\right\}}\ \middle|\ {\mathcal{F}_{k-1}}\right].$
Conditional on $\mathcal{F}_{k-1}$, we can verify that $X_{k},\ldots,X_{n}$ are exchangeable, and so $X_{k}$ has conditional mean $\bar{X}_{\geq k}$ and conditional variance $\sigma^{2}_{X_{\geq k}}$. Since the $X_{i}$’s take values in $[-1,1]$ almost surely, this means $|X_{k}-\bar{X}_{\geq k}|\leq 2$ almost surely. Applying the basic Bernstein inequality (6), since $|w_{k}|\leq\|w\|_{\infty}$, we have
$\mathbb{E}\left[{\exp\left\{\lambda w_{k}(X_{k}-\bar{X}_{\geq k})\right\}}\ \middle|\ {\mathcal{F}_{k-1}}\right]\leq\exp\left\{\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}w_{k}^{2}\sigma^{2}_{X_{\geq k}}\right\}.$
Combining our calculations yields $\mathbb{E}\left[{M_{k}}\ \middle|\ {\mathcal{F}_{k-1}}\right]\leq M_{k-1}$, which completes the proof. ∎
Appendix C Proofs of main results
C.1 Reducing to the case $N=n$
First we verify that, for both Theorem 6 and Theorem 10, it suffices to consider the case $N=n$. To see why, suppose $n<N$. Define
$\tilde{w}=(w_{1},\ldots,w_{n},0,\ldots,0)\in\mathbb{R}^{N}.$
Since $\|\tilde{w}\|_{\infty}=\|w\|_{\infty}$, $\|\tilde{w}\|_{2}=\|w\|_{2}$, and $\sum_{i=1}^{N}\tilde{w}_{i}(X_{i}-\bar{X})=\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})$, by construction, this means that the results of Theorem 6 and Theorem 10 (applied with weight vector $w\in\mathbb{R}^{n}$) are equivalent to the results of these theorems when applied with weight vector $\tilde{w}\in\mathbb{R}^{N}$. Therefore, we only need to prove these theorems for the case $N=n$.
C.2 Proof of Theorem 6
As established in Section C.1, it suffices to consider the case $N=n$. The result of Waudby-Smith and Ramdas [2020, Theorem 3.1] (restated in Proposition 12 above) proves that, for any $v\in\mathbb{R}^{n}$,
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}v_{i}(X_{i}-\bar{X}_{\geq i})-\frac{\lambda^{2}}{2}\|v\|^{2}_{2}\right\}}\right]\leq 1.$
Next, define a matrix $B_{n}\in\mathbb{R}^{n\times n}$ as
$B_{n}=\left(\begin{array}[]{ccccccc}\frac{1}{n}&\frac{1}{n}&\frac{1}{n}&\ldots&\frac{1}{n}&\frac{1}{n}&\frac{1}{n}\\ 0&\frac{1}{n-1}&\frac{1}{n-1}&\ldots&\frac{1}{n-1}&\frac{1}{n-1}&\frac{1}{n-1}\\ 0&0&\frac{1}{n-2}&\ldots&\frac{1}{n-2}&\frac{1}{n-2}&\frac{1}{n-2}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\ldots&\frac{1}{3}&\frac{1}{3}&\frac{1}{3}\\ 0&0&0&\ldots&0&\frac{1}{2}&\frac{1}{2}\\ 0&0&0&\ldots&0&0&1\end{array}\right).$
This matrix provides the following linear operator: for each $i\in[n]$, $(B_{n}X)_{i}=\bar{X}_{\geq i}$, where we write $X=(X_{1},\ldots,X_{n})$. We therefore have
$\sum_{i=1}^{n}v_{i}(X_{i}-\bar{X}_{\geq i})=\sum_{i=1}^{n}v_{i}((\mathbf{I}_{n}-B_{n})X)_{i}=v^{\top}(\mathbf{I}_{n}-B_{n})X.$
Define also $A_{n}\in\mathbb{R}^{n\times n}$ as
$A_{n}=\left(\begin{array}[]{ccccccc}1&-\frac{1}{n-1}&-\frac{1}{n-1}&\ldots&-\frac{1}{n-1}&-\frac{1}{n-1}&-\frac{1}{n-1}\\ 0&1&-\frac{1}{n-2}&\ldots&-\frac{1}{n-2}&-\frac{1}{n-2}&-\frac{1}{n-2}\\ 0&0&1&\ldots&-\frac{1}{n-3}&-\frac{1}{n-3}&-\frac{1}{n-3}\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\vdots\\ 0&0&0&\ldots&1&-\frac{1}{2}&-\frac{1}{2}\\ 0&0&0&\ldots&0&1&-1\\ 0&0&0&\ldots&0&0&0\end{array}\right).$
We can verify that $A_{n}^{\top}$ is the matrix pseudoinverse of $\mathbf{I}_{n}-B_{n}$, and in particular,
$A_{n}^{\top}(\mathbf{I}_{n}-B_{n})=\mathcal{P}^{\perp}_{\mathbf{1}_{n}},$
where $\mathcal{P}^{\perp}_{\mathbf{1}_{n}}$ denotes projection to the subspace orthogonal to $\mathbf{1}_{n}=(1,\ldots,1)\in\mathbb{R}^{n}$.
Now fix any $u\in\mathbb{R}^{n}$ and let $v=A_{n}u$. We then have
$\sum_{i=1}^{n}u_{i}(X_{i}-\bar{X})=u^{\top}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}X=u^{\top}A_{n}^{\top}(\mathbf{I}_{n}-B_{n})X=v^{\top}(\mathbf{I}_{n}-B_{n})X=\sum_{i=1}^{n}v_{i}(X_{i}-\bar{X}_{\geq i}).$
In other words, the MGF bound derived above yields
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}u_{i}(X_{i}-\bar{X})-\frac{\lambda^{2}}{2}\|A_{n}u\|^{2}_{2}\right\}}\right]\leq 1.$ (7)
We pause here to remark that this bound can be viewed as a direct consequence of Waudby-Smith and Ramdas [2020, Theorem 3.1] (restated in Proposition 12)—but if we set $u=w$ so that this result bounds the weighted sum $\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X}_{n})$, this form of the result still does not satisfy aim (c), since $\|A_{n}w\|^{2}_{2}$ is not invariant to permuting the weights $w_{1},\ldots,w_{n}$.
Now we will move towards a result that will treat the weights $w_{i}$ symmetrically. We apply the above bound with $u=(A_{n}^{\top}A_{n})^{\dagger}\Pi w$, where $\Pi\in\{0,1\}^{n\times n}$ is any permutation matrix, while $\dagger$ denotes the matrix pseudoinverse. Note that, since $A_{n}\mathbf{1}_{n}=0$ by construction, this means that $(A_{n}^{\top}A_{n})^{\dagger}\mathbf{1}_{n}=0$ as well—this implies $u^{\top}\mathbf{1}_{n}=0$ and so $\sum_{i=1}^{n}u_{i}(X_{i}-\bar{X})=u^{\top}X=w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}X$. We then have
$\mathbb{E}\left[{\exp\left\{\lambda w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}X-\frac{\lambda^{2}}{2}\|A_{n}(A_{n}^{\top}A_{n})^{\dagger}\Pi w\|^{2}_{2}\right\}}\right]\leq 1,$
for each fixed permutation matrix $\Pi$. Since $X\stackrel{{\scriptstyle\textnormal{d}}}{{=}}\Pi X$ by the exchangeability assumption (where =$\stackrel{{\scriptstyle\textnormal{d}}}{{=}}$ denotes equality in distribution), this implies
$\mathbb{E}\left[{\exp\left\{\lambda w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}\Pi X-\frac{\lambda^{2}}{2}w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}\Pi w\right\}}\right]\leq 1.$
Since this holds for every permutation matrix $\Pi$, we therefore have
$1\geq\frac{1}{n!}\sum_{\Pi}\mathbb{E}\left[{\exp\left\{\lambda w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}\Pi X-\frac{\lambda^{2}}{2}w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}\Pi w\right\}}\right]\\ \geq\mathbb{E}\left[{\exp\left\{\frac{1}{n!}\sum_{\Pi}\left(\lambda w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}\Pi X-\frac{\lambda^{2}}{2}w^{\top}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}\Pi w\right)\right\}}\right],$
where the summation is taken over all $n!$ possible permutation matrices $\Pi$, and where the last step holds by Jensen’s inequality. Moreover, an elementary calculation shows that
$\frac{1}{n!}\sum_{\Pi}\Pi^{\top}(A_{n}^{\top}A_{n})^{\dagger}\Pi=\frac{n-H_{n}}{n-1}\cdot\mathcal{P}^{\perp}_{\mathbf{1}_{n}}=\frac{1}{1+\epsilon_{n}}\cdot\mathcal{P}^{\perp}_{\mathbf{1}_{n}}.$
This means that we have
$1\geq\mathbb{E}\left[{\exp\left\{\frac{1}{1+\epsilon_{n}}\cdot\left(\lambda w^{\top}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}X-\frac{\lambda^{2}}{2}w^{\top}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}w\right)\right\}}\right]\\ =\mathbb{E}\left[{\exp\left\{\frac{1}{1+\epsilon_{n}}\cdot\left(\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})-\frac{\lambda^{2}}{2}w^{\top}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}w\right)\right\}}\right],$
and therefore, since $w^{\top}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}w\leq\|w\|^{2}_{2}$,
$\mathbb{E}\left[{\exp\left\{\frac{1}{1+\epsilon_{n}}\cdot\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right\}}\right]\leq\exp\left\{\frac{1}{1+\epsilon_{n}}\cdot\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\right\}.$
Finally, replacing $\lambda$ with $\lambda\cdot(1+\epsilon_{n})$, we have
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right\}}\right]\leq\exp\left\{(1+\epsilon_{n})\cdot\frac{\lambda^{2}}{2}\|w\|^{2}_{2}\right\},$
as desired.
C.3 Proof of Theorem 8
We will use the bound (7), which was derived as a preliminary step in the proof of Theorem 6. Without loss of generality, assume that we have $w_{1}\geq\ldots\geq w_{n}\geq 0$. Applying this bound with $u=w$, we have
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right\}}\right]\leq\exp\left\{\frac{\lambda^{2}}{2}\|A_{n}w\|^{2}_{2}\right\}.$
For each $i\in[n]$, by definition of $A_{n}$ we have
$\big{(}(A_{n}w)_{i}\big{)}^{2}=(w_{i}-\bar{w}_{\geq i})^{2}\leq w_{i}^{2},$
where the last step holds since, by our assumption on $w$, we have
$w_{i}\geq\bar{w}_{\geq i}\geq 0.$
Therefore, $\|A_{n}w\|^{2}_{2}\leq\|w\|^{2}_{2}$, which completes the proof.
C.4 Proof of Theorem 10
As established in Section C.1, it suffices to consider the case $N=n$. Proposition 13 proves that, for any $v\in\mathbb{R}^{n}$ and any $|\lambda|<\frac{3}{2\|v\|_{\infty}}$,
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}v_{i}(X_{i}-\bar{X}_{\geq i})-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|v\|_{\infty})}\sum_{i=1}^{n}v_{i}^{2}\sigma^{2}_{X_{\geq i}}\right\}}\right]\leq 1.$
As in the proof of Theorem 6, we can rewrite this as
$\mathbb{E}\left[{\exp\left\{\lambda v^{\top}(\mathbf{I}_{n}-B_{n})X-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|v\|_{\infty})}\sum_{i=1}^{n}v_{i}^{2}\sigma^{2}_{X_{\geq i}}\right\}}\right]\leq 1,$
where $B_{n}$ is defined as before. Moreover, for any $i$, we have
$\sigma^{2}_{X_{\geq i}}=\frac{1}{n-i+1}\sum_{j=i}^{n}(X_{j}-\bar{X}_{\geq i})^{2}=\inf_{t\in\mathbb{R}}\left\{\frac{1}{n-i+1}\sum_{j=i}^{n}(X_{j}-t)^{2}\right\}\\ \leq\frac{1}{n-i+1}\sum_{j=i}^{n}(X_{j}-\bar{X})^{2}=\bar{\tilde{X}}_{\geq i}=(B_{n}\tilde{X})_{i},$
where we define a new vector $\tilde{X}$ with entries $\tilde{X}_{i}=(X_{i}-\bar{X})^{2}$. Then
$\sum_{i=1}^{n}v_{i}^{2}\sigma^{2}_{X_{\geq i}}\leq\sum_{i=1}^{n}v_{i}^{2}(B_{n}\tilde{X})_{i}=\tilde{v}^{\top}B_{n}\tilde{X},$
where $\tilde{v}$ is the vector with entries $\tilde{v}_{i}=v_{i}^{2}$. We therefore have
$\mathbb{E}\left[{\exp\left\{\lambda v^{\top}(\mathbf{I}_{n}-B_{n})X-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|v\|_{\infty})}\tilde{v}^{\top}B_{n}\tilde{X}\right\}}\right]\leq 1.$
We now apply the lemma with $v=\Pi w$, where $\Pi\in\{0,1\}^{n\times n}$ is any permutation matrix. Note that $\tilde{v}=\Pi\tilde{w}$ and $\|v\|_{\infty}=\|\Pi w\|_{\infty}=\|w\|_{\infty}$. We then have, for any $|\lambda|<\frac{3}{2\|w\|_{\infty}}$,
$\mathbb{E}\left[{\exp\left\{\lambda w^{\top}\Pi^{\top}(\mathbf{I}_{n}-B_{n})X-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\tilde{w}^{\top}\Pi^{\top}B_{n}\tilde{X}\right\}}\right]\leq 1.$
Since $X\stackrel{{\scriptstyle\textnormal{d}}}{{=}}\Pi X$, we can replace $X$ with $\Pi X$ in the expression above to obtain
$\mathbb{E}\left[{\exp\left\{\lambda w^{\top}\Pi^{\top}(\mathbf{I}_{n}-B_{n})\Pi X-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\tilde{w}^{\top}\Pi^{\top}B_{n}\Pi\tilde{X}\right\}}\right]\leq 1,$
since $\widetilde{\Pi X}=\Pi\tilde{X}$ (because $(\widetilde{\Pi X})_{i}=((\Pi X)_{i}-\overline{(\Pi X)})^{2}=((\Pi X)_{i}-\bar{X})^{2}=(\Pi\tilde{X})_{i}$). Since this holds for every permutation matrix $\Pi$, we therefore have
$1\geq\frac{1}{n!}\sum_{\Pi}\mathbb{E}\left[{\exp\left\{\lambda w^{\top}\Pi^{\top}(\mathbf{I}_{n}-B_{n})\Pi X-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\tilde{w}^{\top}\Pi^{\top}B_{n}\Pi\tilde{X}\right\}}\right]\\ \geq\mathbb{E}\left[{\exp\left\{\frac{1}{n!}\sum_{\Pi}\left(\lambda w^{\top}\Pi^{\top}(\mathbf{I}_{n}-B_{n})\Pi X-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\tilde{w}^{\top}\Pi^{\top}B_{n}\Pi\tilde{X}\right)\right\}}\right],$
where the summation is taken over all $n!$ possible permutation matrices $\Pi$, and where the last step holds by Jensen’s inequality. A straightforward calculation verifies that
$\frac{1}{n!}\sum_{\Pi}\Pi^{\top}B_{n}\Pi=\mathbf{I}_{n}-\frac{n-H_{n}}{n-1}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}=\mathbf{I}_{n}-\frac{1}{1+\epsilon_{n}}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}=\frac{1}{1+\epsilon_{n}}\mathcal{P}_{\mathbf{1}_{n}}+\frac{\epsilon_{n}}{1+\epsilon_{n}}\mathbf{I}_{n},$
and so we have
$\mathbb{E}\left[{\exp\left\{\lambda w^{\top}\left(\frac{1}{1+\epsilon_{n}}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}\right)X-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\tilde{w}^{\top}\left(\frac{1}{1+\epsilon_{n}}\mathcal{P}_{\mathbf{1}_{n}}+\frac{\epsilon_{n}}{1+\epsilon_{n}}\mathbf{I}_{n}\right)\tilde{X}\right\}}\right]\leq 1.$
We can also calculate $w^{\top}\mathcal{P}^{\perp}_{\mathbf{1}_{n}}X=\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})$, and also,
$\tilde{w}^{\top}\mathcal{P}_{\mathbf{1}_{n}}\tilde{X}=\sum_{i=1}^{n}\tilde{w}_{i}\cdot\frac{\sum_{i=1}^{n}\tilde{X}_{i}}{n}=\|w\|^{2}_{2}\sigma^{2}_{X}$
and
$\tilde{w}^{\top}\mathbf{I}_{n}\tilde{X}=\sum_{i=1}^{n}\tilde{w}_{i}\tilde{X}_{i}=\sum_{i=1}^{n}w_{i}^{2}(X_{i}-\bar{X})^{2}\leq 4\|w\|^{2}_{2}.$
Plugging in these calculations, then,
$\mathbb{E}\left[{\exp\left\{\frac{\lambda}{1+\epsilon_{n}}\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})-\frac{\lambda^{2}}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty})}\left(\frac{\|w\|^{2}_{2}\sigma^{2}_{X}}{1+\epsilon_{n}}+\frac{\epsilon_{n}\cdot 4\|w\|^{2}_{2}}{1+\epsilon_{n}}\right)\right\}}\right]\leq 1.$
Replacing $\lambda$ with $\lambda(1+\epsilon_{n})$ (and assuming now that $|\lambda|<\frac{3}{2\|w\|_{\infty}(1+\epsilon_{n})}$) yields
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})-\frac{\lambda^{2}(1+\epsilon_{n})}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty}(1+\epsilon_{n}))}\|w\|^{2}_{2}(\sigma^{2}_{X}+4\epsilon_{n})\right\}}\right]\leq 1.$
This completes the proof of the theorem.
C.5 Proofs of Corollaries 79, and 11
For the Hoeffding bounds, Corollaries 7 and 9 follow immediately from Theorems 6 and 8, respectively, by applying the fact (1).
For the Bernstein bound, we need an additional step to verify Corollary 11. First, observe that $X=(X_{1},\dots,X_{N})$ is exchangeable conditional on $\tilde{\sigma}^{2}_{X,N}$ (because $\tilde{\sigma}^{2}_{X,N}$ is a symmetric function of $X$). Applying Theorem 10 to this conditional distribution, then, we obtain
$\mathbb{E}\left[{\exp\left\{\lambda\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\right\}}\ \middle|\ {\tilde{\sigma}^{2}_{X,N}}\right]\leq\exp\left\{\frac{\lambda^{2}(1+\epsilon_{N})}{2(1-\frac{2|\lambda|}{3}\|w\|_{\infty}(1+\epsilon_{N}))}\cdot\tilde{\sigma}^{2}_{X,N}\|w\|^{2}_{2}\right\}$
for any $|\lambda|<\frac{3}{2\|w\|_{\infty}(1+\epsilon_{N})}$. By the fact (2) applied with $a=(1+\epsilon_{N})\tilde{\sigma}^{2}_{X,N}\|w\|^{2}_{2}$ and $b=\frac{2}{3}\|w\|_{\infty}(1+\epsilon_{N})$, we then have the following bound on the conditional tail probability:
$\mathbb{P}\left\{{\sum_{i=1}^{n}w_{i}(X_{i}-\bar{X})\geq\tilde{\sigma}_{X,N}\|w\|_{2}\sqrt{2(1+\epsilon_{N})\log(1/\delta)}+\frac{2}{3}\|w\|_{\infty}(1+\epsilon_{N})\log(1/\delta)}\ \middle|\ {\tilde{\sigma}^{2}_{X,N}}\right\}\leq\delta.$
After marginalizing over $\tilde{\sigma}^{2}_{X,N}$, this yields the desired bound.